OpenStudy (anonymous):

Calculate: 20 + 3(5 − 1)

OpenStudy (anonymous):

5-1 = 4 3(5-1) = 3*4 = 12 20 + 12 = 32

OpenStudy (anonymous):

that is the correct answer i just double checked it

OpenStudy (anonymous):

thankss!